OpenStudy (anonymous):

If \(\alpha\) and \(\beta\) are the roots of the equation \(x^2 px + q =0\) and \(x^{2n} + p^n x^n + q^n =0\) and if \(\left(\frac{\alpha}{\beta},\frac{\beta}{\alpha}\right)\) are the roots of \(x^n + 1 + (x + 1)^n = 0\), then \(n\) must be a. Odd Integer c. Rational but not Integer. b. Even Integer d. None of these.

OpenStudy (anonymous):

\(x^2 + px +q = 0\) not \(x^2 px + q = 0\)

OpenStudy (jamesj):

That last equation is right?

OpenStudy (anonymous):

yeah

OpenStudy (jamesj):

Well, n can't be even then. Now ...

OpenStudy (anonymous):

Wow, but How?

OpenStudy (jamesj):

From that last equation. If n is even, every quantity (..stuff...)^n is non-negative and we have +1 as well, all of it = 0. That's a contradiction.

OpenStudy (jamesj):

[ Formally. If n = 2k, for some integer k, then \[ 0 = x^{2k} + 1 + (x+1)^{2k} = (x^2)^k + 1 + ((x+1)^2)^k \geq 0 + 1 + 0 = 1 \] contradiction.] Now, from the first equation \[ \alpha + \beta = -p --- (1)\] \[ \alpha . \beta = q ---(2) \] From the second ... this is bit more subtle, think of it as a quadratic in x^n... \[ \alpha^n + \beta^n = -p^n --- (3) \] \[ \alpha^n . \beta^n = q^n --- (4) \] (1) => (4), so that's nothing new. But clearly (1) and (3) are different. Looking now at (1) and taking the nth power of both sides and using n as an odd integer, \[ (\alpha + \beta)^n = (-p)^n = -p^n --- (5) \] From the 3rd equation of the problem \[ -[ (\alpha/\beta)^n + 1] = (\alpha/\beta+1)^n --- (6) \] In equation (5), divide both sides by beta^n, \[ (\alpha/\beta+1)^n = -p^n/\beta^n \] By equation (6), \[ -p^n/\beta^n = -[ (\alpha/\beta)^n + 1] \implies p^n = (\alpha + \beta)^n\] consistent with equation (1). Therefore, in short, the system of equations are consistent for n being an odd integer.

OpenStudy (jamesj):

**correction, in that last deduction it should be \[ ... \implies p^n = \alpha^n + \beta^n \] consistent with equation (3).

OpenStudy (jamesj):

Where's this question from? It looks like it could be from the Math GRE.

OpenStudy (anonymous):

Thanks JamesJ I don't know, my teacher gave it to me.

OpenStudy (jamesj):

I'd like to point out to the purists that while I have shown option A isn't an option, I haven't shown C and D cannot be options. But I have shown B is a possibility and I'm now leaning on the fact that only one option is correct. I suspect it wouldn't be too hard to show that C and D lead to a contradiction; I can see how it would pivot around the one small result I used in equation (5): the treatment of (-p)^n. **One other correction, just after equation 4: "(2) [*not (1)] => (4), so that's nothing new."